# On the Growth of Mistakes in Differentially Private Online Learning: A Lower Bound Perspective

### v1

@@ 4.2 Examples of $\operatorname{\beta-concentrated}$ online learners for $\mathrm{Point}_{N}$
For $K\leqslant N\in\mathbb{N}_{+}$, we define
$\mathrm{Point}^{K}_{N}\coloneqq\left\{f\in\left\{0,1\right\}^{\mathcal{X}},\text{ such that }\left\lvert\operatorname{supp}f\right\rvert\leqslant K\right\},$ (4.10)
[…]
@@
We refer reader to Appendix A for the proof. We remark that the SOA for $\mathrm{Point}_{N}$, which is an improper learner, uses $\mathrm{Point}_{N}^{1}$eimel et al. [6] also uses improper learners consisting of union of points for privately learning $\mathrm{Point}_{N}$ in the pure DP-PAC model; their hypothesis class is a subset of the Multi-Point class.333 See Section 4.1 in Beimel et al. [6] to see the conditions that the hypothesis class needs to satisfy.
@@ 5.1 Connection to DP under continual observation
Several works (see Chan et al. [10]) have identified counters with $\alpha=O\left(\left({\log t}\right)^{1.5}/\varepsilon\right)$ which immediately implies a mistake bound that also scales as $\left({\log t}\right)^{1.5}/\varepsilon$ using Section 5.1. Ignoring [^4] We do not optimise this as this dependence is not the focus of this work, for this argument consider $N=O\left(1\right)$. However, we believe it can be reduced using a DP continual algorithm for MAXSUM; see Jain et al. [21] the dependence on $N$, this almost matches the lower bound proposed in Theorem 1. On the other hand, Dwork et al. [15] have shown a lower bound of $\Omega\left({\frac{\log T}{\varepsilon}}\right)$ for $\alpha$ for any pure DP-continual counter. However it is not clear how to convert the lower bound for DP-continual learning to DP online learning. This is because 1) Equation 5.1 asks for a uniformly (across $t$) accurate counter whereas the mistake bound in Equation 2.1 is a global measure and 2) this lower bound is for pure DP whereas our setting is approximate DP. We leave this question for the future work.
@@ 5.2 Beyond concentration assumption
Proof is provided in Appendix C and is similar to the proof of Theorem 1, but requires a more delicate construction.
@@ 5.3 Pure differentially private online learners
For $(\varepsilon,\delta)$-DP online algorithms, there exists an upper bound provided by Golowich and Livni [19]. However, to the best of our knowledge, not much is known about $\varepsilon$-DP online algorithms. One way to obtain such algorithm is by leveraging existing results from the continual observation literature as done in Section 5.1. Under the same assumptions as in Section 5.1, using basic composition instead of advanced composition in the last step results in an $\varepsilon^{\prime}$ scaling as $N\varepsilon$. However, this only works for $\mathrm{Point}_{N}$ and not for more general classes. For arbitrary finite hypothesis classes, it is possible to use a DP continual counter, similar to above to obtain a mistake bound that also scales with the size of the class. However, the question of whether generic learners can be designed remains open. Another interesting open question is whether the dependence on the size of the class e.g., $N$ for $\mathrm{Point}_{N}$, is necessary.
@@ 6 Acknowledgement
DD is supported by ETH AI Center doctoral fellowship and ETH Foundations of Data Science initiative. AS and DD thank Yevgeny Seldin for hosting them in University of Copenhagen which helped in the discussion of this work. DD thanks Afonso Bandeira and Petar Nizić-Nikolac for fruitful discussions during the course of the project. AS thanks Amir Yehudayoff for helpful discussions.
Appendix A Proof of Lemma 1
$\sum_{x\in[N]}\operatorname*{\mathbb{P}}(\hat{f}_{t}(x)=1)$ $=\sum_{x\in[N]}\sum_{\begin{subarray}{c}f\in\mathrm{Point}^{K}_{N}\\ f(x)=1\end{subarray}}\operatorname*{\mathbb{P}}(\hat{f}_{t}=f)$ (A.1)
$=\sum_{f\in\mathrm{Point}^{K}_{N}}\sum_{\begin{subarray}{c}x\in[N]\\ f(x)=1\end{subarray}}\operatorname*{\mathbb{P}}(\hat{f}_{t}=f)$
$\leqslant K\sum_{f\in\mathrm{Point}^{K}_{N}}\operatorname*{\mathbb{P}}(\hat{f}_{t}=f)=K,$
$\operatorname*{\mathbb{P}}(\exists t,\text{ such that }\hat{f}_{t}(x_{\mathrm{dif}})=1)\leqslant\sum_{t=1}^{T}\operatorname*{\mathbb{P}}(\hat{f}_{t}(x_{\mathrm{dif}})=1)\leqslant\beta,$ (A.2)
which implies that $\mathcal{A}$ is $\operatorname{\beta-concentrated}$ and corresponding distinguishing tuple is $(f_{1},{\mathbb{I}}\left\{{\cdot=x_{\mathrm{dif}}}\right\},x_{\mathrm{eq}},x_{\mathrm{dif}})$.
In Lemma 1 we assumed that $K$ is a constant. Note that as long as $K=o(N)$, by taking $N$ large enough one can obtain the same result.
Appendix B Proof of Section 5.1
@@ Appendix A Proof of Lemma 1
Appendix C Lower bound for a non-concentrated algorithm
@@
$\operatorname*{\mathbb{P}}(\mathcal{A}(\tau_{i})_{t}=f^{\ast})$ $={\mathbb{P}}_{i}(\mathcal{A}(\tau_{i})_{t}=f^{\ast})\operatorname*{\mathbb{P}}(G_{i})+\operatorname*{\mathbb{P}}(\mathcal{A}(\tau_{i})_{t}=f^{\ast}\mid\overline{G_{i}})\operatorname*{\mathbb{P}}(\overline{G_{i}})$ (C.1)
@@
${\mathbb{P}}_{i}(\mathcal{A}(\tau_{i})\text{ outputs only }f^{\ast}\text{ starting from }r_{i})$ (C.2)
@@
${\mathbb{P}}_{i}(\mathcal{A}(\tau_{i})\text{ outputs only }f^{\ast}\text{ on }[lm_{i},rm_{i}])$ (C.3)
@@
${\mathbb{P}}_{i}(\mathcal{A}(\tau_{i+1})\text{ outputs only }f^{\ast}\text{ starting from }r_{i})$ (C.4)
@@
$\operatorname*{\mathbb{P}}(\mathcal{A}(\tau_{i+1})\text{ outputs only }f^{\ast}\text{ starting from }r_{i}\textbf{ and }\text{ outputs only }f^{\ast}\text{ on }[lm_{i},rm_{i}])$ (C.5)
@@
$\operatorname*{\mathbb{P}}(\mathcal{A}(\tau_{i+1}\text{ outputs only }f^{\ast}\text{ starting from }r_{i}\textbf{ and }\text{does not only output }f^{\ast}\text{ on }[lm_{i},rm_{i}])$ (C.6)
@@
$\operatorname*{\mathbb{P}}(\mathcal{A}(\tau_{i+1})\text{ outputs only }f^{\ast}\text{ starting from }lm_{i})\leqslant\operatorname*{\mathbb{P}}(\mathcal{A}(\tau_{i+1})\text{ outputs all ones on }[lm_{i},rm_{i}]).$ (C.7)
@@
Proof of Section 5.2 resembles the proof of Theorem 1 with several important differences. The main difficulty comes from the fact that by just looking at the output of the algorithm, it is impossible to say with certainty, whether it ‘fired or not. And we need to do this, in order to bound the marginal probability of outputting $f^{\ast}$ (which is $1/2$ if $\mathcal{A}$ did not fire and $1$ if it fired), see Equation C.1.
To overcome this issue, instead of splitting the segment into two parts and reiterating in either left or right, we split into three parts and reiterate in the first or the third. Second segment plays a special role, as to ’detect’ whether algorithm fired or not. From the properties of the algorithm, if the full segment is equal to all $f^{\ast}$, we can confidently say that algorithm fired before this segment. Event $G_{i}$ corresponds to a ’good’ event, meaning that $\mathcal{A}$ did not look like it ’fired’ even though it did not.
@@
Appendix D Extending finite lower bound to other values of $\varepsilon$
@@
$\min\left({\operatorname*{\mathbb{P}}\left[{Q_{1}(\bm{c}^{(i+1)})}\right],\operatorname*{\mathbb{P}}\left[{Q_{2}(\bm{c}^{(i+1)})}\right]}\right)\leqslant\frac{1}{2}\left({\exp(\varepsilon_{0})\operatorname*{\mathbb{P}}\left[{Q(\bm{c}^{(i)})}\right]+A\delta}\right),$ (D.1)
@@
$\min\left({\operatorname*{\mathbb{P}}\left[{Q_{1}(\bm{c}^{(i+1)})}\right],\ldots,\operatorname*{\mathbb{P}}\left[{Q_{s}(\bm{c}^{(i+1)})}\right]}\right)\leqslant\frac{1}{s}\left({\exp(\varepsilon)\operatorname*{\mathbb{P}}\left[{Q(\bm{c}^{(i)})}\right]+\delta}\right).$ (D.2)

### v3

@@ 4.2 Examples of $\operatorname{\beta-concentrated}$ online learners for $\mathrm{Point}_{N}$
For $1\leqslant K\leqslant N\in\mathbb{N}_{+}$, we define $\mathrm{Point}^{K}_{N}$ to be the class of functions that are equal to 1 on at most $K$ points.
@@
$\sum_{x\in[N]}\operatorname*{\mathbb{P}}(\hat{f}_{t}(x)=1)=\sum_{x\in[N]}\sum_{\begin{subarray}{c}f\in\mathrm{Point}^{K}_{N}\\ f(x)=1\end{subarray}}\operatorname*{\mathbb{P}}(\hat{f}_{t}=f)$ (4.11)
$\quad=\sum_{f\in\mathrm{Point}^{K}_{N}}\sum_{\begin{subarray}{c}x\in[N]\\ f(x)=1\end{subarray}}\operatorname*{\mathbb{P}}(\hat{f}_{t}=f)\leqslant K\sum_{f\in\mathrm{Point}^{K}_{N}}\operatorname*{\mathbb{P}}(\hat{f}_{t}=f)=K,$
$\operatorname*{\mathbb{P}}(\exists t,\text{ such that }\hat{f}_{t}(x_{\mathrm{dif}})=1)\leqslant\sum_{t=1}^{T}\operatorname*{\mathbb{P}}(\hat{f}_{t}(x_{\mathrm{dif}})=1)\leqslant\beta.$ (4.12)
Thus, $\mathcal{A}$ is $\operatorname{\beta-concentrated}$ and the distinguishing tuple is $(f_{1},{\mathbb{I}}\left\{{\cdot=x_{\mathrm{dif}}}\right\},x_{\mathrm{eq}},x_{\mathrm{dif}})$.
In Lemma 1 we assumed that $K$ is a constant. Note that as long as $K=o(N)$, by taking $N$ large enough one can obtain the same result. We also remark that the SOA for $\mathrm{Point}_{N}$, which is an improper learner, uses $\mathrm{Point}_{N}^{1}$ as the output hypothesis class. Beimel et al. [6] also uses improper learners consisting of union of points for privately learning $\mathrm{Point}_{N}$ in the pure DP-PAC model; their hypothesis class is a subset of the Multi-Point class.333 See Section 4.1 in Beimel et al. [6] to see the conditions that the hypothesis class needs to satisfy.
@@
Independent work In a concurrent and independent work, Cohen et al. [12] established that learners that are not concentrated must nevertheless suffer large expected regret for the specific case of $\mathrm{Point}_{T}$. They construct an adversary which strongly relies on the large size of the function class. In particular, generalizing for a smaller function classes, e.g. $\mathrm{Point}_{3}$ remains an interesting open question. We obtain a partial progress in this direction, see Section 5.2.
@@ 5.1 Connection to DP under continual observation
Several works (see Chan et al. [10]) have identified counters with $\alpha=O\left(\left({\log T}\right)^{1.5}/\varepsilon\right)$ which immediately implies a mistake bound that also scales as $\operatorname{poly}\log\left({T}\right)/\varepsilon$ using Section 5.1. Ignoring [^4] We do not optimise this as this dependence is not the focus of this work, for this argument consider $N=O\left(1\right)$. However, we believe it can be reduced using a DP continual algorithm for MAXSUM; see Jain et al. [23] the dependence on $N$, this almost matches the lower bound proposed in Theorem 1. On the other hand, Dwork et al. [16] have shown a lower bound of $\Omega\left({\frac{\log T}{\varepsilon}}\right)$ for $\alpha$ for any pure DP-continual counter. However it is not clear how to convert the lower bound for DP-continual learning to DP online learning. This is because 1) Equation 5.1 asks for a uniformly (across $t$) accurate counter whereas the mistake bound in Equation 2.1 is a global measure and 2) this lower bound is for pure DP whereas our setting is approximate DP. We leave this question for the future work.
@@ 5.2 Beyond concentration assumption
Proof is provided in Appendix B and is similar to the proof of Theorem 1, but requires a more delicate construction.
@@ 5.3 Pure differentially private online learners
For $(\varepsilon,\delta)$-DP online algorithms, there exists an upper bound provided by Golowich and Livni [21]. However, to the best of our knowledge, not much is known about $\varepsilon$-DP online algorithms. One way to obtain such algorithm is by leveraging existing results from the continual observation literature as done in Section 5.1. Under the same assumptions as in Section 5.1, using basic composition instead of advanced composition in the last step results in an $\varepsilon^{\prime}$ scaling as $N\varepsilon$. However, this only works for $\mathrm{Point}_{N}$ and not for more general classes. For arbitrary finite hypothesis classes, it is possible to use a DP continual counter, similar to above to obtain a mistake bound that also scales with the size of the class. We also remark that the $\mathsf{AboveThreshold}$ algorithm could be used to design learners for certain function classes. However, the question of whether generic learners can be designed remains open. Another interesting open question is whether the dependence on the size of the class e.g., $N$ for $\mathrm{Point}_{N}$, is necessary.
5.4 Open problems
This work relies on assumptions about properties (Concentrated Assumption in Definition 7 or uniform firing Assumption in Section 5.2) of the learning algorithms to show a lower bound for any hypothesis class. On the other hand, Cohen et al. [12] do not need any assumption on their algorithm but their result only holds for large $\mathrm{Point}$ classes (and it cannot immediately be transferred to small $\mathrm{Point}$ classes, e.g. $\mathrm{Point}_{3}$). To fully settle the problem of lower bounds for online DP learners, the limitations of both this work and that of Cohen et al. [12] need to be addressed. We propose the following conjecture which we believe is true.
Conjecture 6 (Lower bound for Approximate DP)
For any $\varepsilon,\delta>0$ and any $(\varepsilon,\delta)$-DP learner $\mathcal{A}$ of $\mathrm{Point}_{3}$, there exists an adversary $\mathcal{B}$, such that $\operatorname*{\mathbb{E}}\left[{M_{\mathcal{A}}}\right]=\Omega\left({\min\left({\frac{\log T/\delta}{\varepsilon},1/\delta}\right)}\right)$.
A straightforward implication of our main result is that the $\log T$ lower bound also holds for pure DP online learners under the same assumptions on the algorithm. In Section 5.1, we showed a generic reduction from DP continual counters to DP online learners for $\mathrm{Point}_{N}$ with regret $\operatorname{poly}\log\left({T}\right)$. This reduction can be extended to pure DP online learner by using basic composition in Section 5.1, with an additional cost of $\sqrt{N}$. In Section 5.4, we raise the question whether for small $N$, the dependence on $T$ can be lowered to $\log T$ and shown to be tight.
Conjecture 7 (Upper and lower bounds for Pure DP)
For any $\varepsilon>0$,
  1. 1.
    Upper bound: there exists $\varepsilon$-DP learner of $\mathrm{Point}_{3}$, s.t. for any adversary, $\operatorname*{\mathbb{E}}\left[{M}\right]=O\left({\frac{\log T}{\varepsilon}}\right)$.
  2. 2.
    Lower bound: for any $\varepsilon$-DP learner of $\mathrm{Point}_{3}$, there exists an adversary, s.t. $\operatorname*{\mathbb{E}}\left[{M}\right]=\Omega\left({\frac{\log T}{\varepsilon}}\right)$.
Note that the lower bound part of Section 5.4 follows from Section 5.4, but can also be viewed through connection to the continual observation model. In the latter regime, an $\Omega(\log T)$ lower bound was shown in Dwork et al. [18].
@@ 6 Acknowledgement
DD is supported by ETH AI Center doctoral fellowship and ETH Foundations of Data Science initiative. AS and DD thank Yevgeny Seldin for hosting them in University of Copenhagen which helped in the discussion of this work and Amir Yehudayoff for helpful discussions. DD also thanks Afonso Bandeira and Petar Nizić-Nikolac for fruitful discussions during the course of the project. In addition, authors would like to thank Xiaoyu Wang for pointing out a mistake in the proof of Theorem 1 in the earlier version which has now been fixed.
Appendix A Proof of Section 5.1
@@
Appendix B Lower bound for a non-concentrated algorithm
@@
$\operatorname*{\mathbb{P}}(\mathcal{A}(\tau_{i})_{t}=f^{\ast})$ $={\mathbb{P}}_{i}(\mathcal{A}(\tau_{i})_{t}=f^{\ast})\operatorname*{\mathbb{P}}(G_{i})+\operatorname*{\mathbb{P}}(\mathcal{A}(\tau_{i})_{t}=f^{\ast}\mid\overline{G_{i}})\operatorname*{\mathbb{P}}(\overline{G_{i}})$ (B.1)
@@
${\mathbb{P}}_{i}(\mathcal{A}(\tau_{i})\text{ outputs only }f^{\ast}\text{ starting from }r_{i})$ (B.2)
@@
${\mathbb{P}}_{i}(\mathcal{A}(\tau_{i})\text{ outputs only }f^{\ast}\text{ on }[lm_{i},rm_{i}])$ (B.3)
@@
${\mathbb{P}}_{i}(\mathcal{A}(\tau_{i+1})\text{ outputs only }f^{\ast}\text{ starting from }r_{i})$ (B.4)
@@
$\operatorname*{\mathbb{P}}(\mathcal{A}(\tau_{i+1})\text{ outputs only }f^{\ast}\text{ starting from }r_{i}\textbf{ and }\text{ outputs only }f^{\ast}\text{ on }[lm_{i},rm_{i}])$ (B.5)
@@
$\operatorname*{\mathbb{P}}(\mathcal{A}(\tau_{i+1}\text{ outputs only }f^{\ast}\text{ starting from }r_{i}\textbf{ and }\text{does not only output }f^{\ast}\text{ on }[lm_{i},rm_{i}])$ (B.6)
@@
$\operatorname*{\mathbb{P}}(\mathcal{A}(\tau_{i+1})\text{ outputs only }f^{\ast}\text{ starting from }lm_{i})\leqslant\operatorname*{\mathbb{P}}(\mathcal{A}(\tau_{i+1})\text{ outputs all ones on }[lm_{i},rm_{i}]).$ (B.7)
@@
Proof of Section 5.2 resembles the proof of Theorem 1 with several important differences. The main difficulty comes from the fact that by just looking at the output of the algorithm, it is impossible to say with certainty, whether it ‘fired’ or not. And we need to do this, in order to bound the marginal probability of outputting $f^{\ast}$ (which is $1/2$ if $\mathcal{A}$ did not fire and $1$ if it fired), see Equation B.1.
To overcome this issue, instead of splitting the segment into two parts and reiterating in either left or right, we split into three parts and reiterate in the first or the third. Second segment plays a special role, to ’detect’ whether algorithm fired or not. From the properties of the algorithm, if the full segment is equal to all $f^{\ast}$, we can confidently say that algorithm fired before this segment. Event $G_{i}$ corresponds to a ’good’ event, meaning that $\mathcal{A}$ did not fire and did not look like it fired, i.e., in the middle segment there exist one output not equal to $f^{\ast}$.
@@
Appendix C Extending finite lower bound to other values of $\varepsilon$
@@
$\min\left({\operatorname*{\mathbb{P}}\left[{Q_{1}(\bm{c}^{(i+1)})}\right],\operatorname*{\mathbb{P}}\left[{Q_{2}(\bm{c}^{(i+1)})}\right]}\right)\leqslant\frac{1}{2}\left({\exp(\varepsilon_{0})\operatorname*{\mathbb{P}}\left[{Q(\bm{c}^{(i)})}\right]+A\delta}\right),$ (C.1)
@@
$\min\left({\operatorname*{\mathbb{P}}\left[{Q_{1}(\bm{c}^{(i+1)})}\right],\ldots,\operatorname*{\mathbb{P}}\left[{Q_{s}(\bm{c}^{(i+1)})}\right]}\right)\leqslant\frac{1}{s}\left({\exp(\varepsilon)\operatorname*{\mathbb{P}}\left[{Q(\bm{c}^{(i)})}\right]+\delta}\right).$ (C.2)